• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Bosted

    Log entry time 16:08:40 on November23,2008

    Entry number 250467

    keyword=Day shift summary

    Shift started with smooth data taking

    08:30 Due to some problems with the analyzer, we weren't sure if there
    was good BigByte data for runs 3835-37. So we restarted CODA again. It
    turns out these runs are fine after all.

    11:00 MCC took the beam for five minutes for orbit studies

    12:00 a few special runs were made with no pedestal suppression.

    Smooth data taking until end of the shift

    run Coulombs good?
    3837 0.007 yes
    3838 0. ignore
    3839 0.035 good
    3840 0.035 good
    3841 0.035 good
    3842 0.001 test: no ped suppression
    3843 0.005 test, ped. supression off with right prescales
    3844 0.035 good
    3845 0.035 good
    3846 - short, ignore (happex daq off)
    3847 - short - ignore (happex daq not working)
    3848 0.036 good